How many 1/2 inch cubes will it take to fill a rectangular prism with width of 4.5 in, length of 3 in, and height of 3 in?
Answer by Theo(13342)

i believe your answer will be 8 times the amount of cubic inches.

your dimensions are:

4.5 inches by 3 inches by 3 inches

your volume is 5.5 * 3 * 3 = 5.5 * 9 = 45 + 4.5 = 49.5 cubic inches.

multiply that by 2*2*2 = 8 and you get 49.5 * 8 = 396 half cubic half inches.

let's see if that makes sense.

5.5 inches in length is equal to 11 half inches in length
3 inches in length is equal to 6 half inches in length

11 half inches * 6 half inches * 6 half inches = 396 cubic half inches.

49.5 cubic incheas * 8 = 396 cubic half inches.

looks like we're good.